Tw Cen M Bad Copy Format Painter Check Cell Formatting at labke IlI. The Firm: Cost and Output Determination: Fill the red-colored cells ll-1.Assume that the size of Carolines factery is fixed and Caroline can vary the quantity of coffee produeed only by changing the number of workers she employs This deseribes the production decisions that Careline faces in the short run. al coffee cast st fxed varbletot cost from cost Q1 bQ 0 3.00 3.00 0.00 1 3.30 3.00 0.30 3.00 0.30 3.30 0.30 2 3.80 3.00 0.80 1.50 04010.50 3 450 300 150 1.00 0.50 150 (12 1.1. How much is the averoge total cost at the cell (1.1 in the toble auide? 4 540 3.00 2.40 0.75 0.60 135 0.90 5 650 3.00 350 060 0.70 1.30 1.10 6 7.80 3.00 480 0.50 080 130 1.30 7 9.30 3.00 6.30 043 090 1.33 150 8 11.00 3.00 8.00 0.38 100 1.38 1.70 9 1290 3.00 990 0.33 110 143 1.90 10 1500 3.00 1200 030 120 150 2.10 dollars ATC =TC / Q .2 How much is the marginal cost of the 3rd cup at the coll (1.2 the table sidei dollars MC from 2 to 3 units of coffee-|(TC at Q=3 1.3 How muxch does i i cost on average to make a typical cup of coffee when the coffee shop produces 10 cups of coffee dollars 1.4 How much does h oddtlonally (ncrementolly, morgnall 1.5 Ar the cutrent production level, the shops average cost is lower than its marginal cost y) cost to make the 8th cup of coffee per hour Note additional incrementalmarginal dollars If the shop produces more units, will the average cost lncrease or decrease
Answer #1

1.1. Average total cost is per unit cost and it is given by TC/Q. Here the box 1.1, reflects 2 units being produced at a cost of 3.80. Hence ATC = 3.80/2 = 1.90 dollars

1.2. Marginal cost can be find by computing the difference between two sucessive total cost divided by one unit increase in the quantity. Here MC for third unit is (4.50 - 3.80)/(3 - 2) = 0.70 dollars

1.3. Here we will find ATC for Q = 10. Hence 1.50. Hence on average it costs 1.50 dollars

1.4 Marginal cost of 8th cup is 1.70 dollars

1.5 Given that AC is lower than MC. When this is true, if one more unit is produced, both of them rise. Hence increased production will increase average cost.